# Three Dimensional Topological Field Theories and Nahm Sum Formulas

### v1

@@ 3.3 Simple lines
Suppose that an ${\cal N}=2$ gauge theory ${\cal T}[K,\{{\cal O}_{I}\}]$ flows to an ${\cal N}=4$ rank-zero SCFT. Then there exists a distinguished set of line operators $\{L_{1}^{\text{UV}},\cdots L_{d}^{\text{UV}}\}$ in the UV gauge theory which flow to simple lines (i.e. anyons) in the IR topologically twisted field theory. In this section, we focus on the case where UV line operators take the form of supersymmetric Wilson loops $W_{\vec{Q}}$ with charge $\vec{Q}=(Q_{1},\ldots,Q_{r})$, and denote by $W^{\rm sim}_{A/B}$ the collection of loops which map to simple lines. The two sets $W^{\rm sim}_{A}$ and $W^{\rm sim}_{B}$ always contain the trivial Wilson line, $1=W_{\vec{Q}=\vec{0}}$, which corresponds to the trivial simple line in the twisted TFT. Below we will list several necessary conditions for a UV operator $W_{\vec{Q}}$ to become a simple object in the IR.
@@
In order to test whether a UV loop $W_{\vec{Q}}$ flows to a simple line in the IR A-twisted TFT we scan over integer charge vectors $\vec{Q}$ and search for $W_{\vec{Q}}$ which satisfy [^9] More precisely, the conditions for $W_{\vec{Q}}$ to become a simple object, up to a factor of $U(1)_{R}$ flavor Wilson loop which contributes $(-q^{\pm 1/2})^{Q_{R}\in\mathbb{Z}}$ to $\langle W_{\vec{Q}}^{\pm}\rangle_{\rm sci}$. These contributions cancel in cancel in $\langle W^{+}_{\vec{Q}}W^{-}_{\vec{Q}}\rangle_{\rm sci}$.
@@
The first condition follows from the fact that the $S^{2}\times S^{1}$ partition function for a semi-simple TFT with a simple line insertion either vanishes or is equal to 1 if we insert the identity line. The second condition follows from the fact that the fusion between a simple object $\mathbf{L}$ and its orientation reversal $\overline{\mathbf{L}}$ always contains a single $1$ (trivial object), i.e.,
@@
More concretely, we have
$\begin{split}&{\cal Z}[S^{2}\times S^{1}\textrm{ with two simple objects }\mathbf{L}_{\alpha}\textrm{ and }\mathbf{L}_{\beta}\textrm{ along the}S^{1}]\\ =&~{}\sum_{\gamma}S_{0\gamma}^{2}\frac{S_{\alpha\gamma}}{S_{0\gamma}}\frac{S_{\beta\gamma}}{S_{0\gamma}}=(S^{2})_{\alpha\beta}\\ =&~{}\begin{cases}1,\quad\textrm{if }\mathbf{L}_{\alpha}=\overline{\mathbf{L}}_{\beta}\\ 0,\quad\textrm{otherwise }\end{cases}\end{split}$ (35)
Here $S_{\alpha\beta}$ is the modular S-matrix and $(S^{2})_{\alpha\beta}=C_{\alpha\beta}$ is the charge-conjugation matrix. If a UV Wilson loop maps to a sum of multiple non-trivial simple objects in the IR, it satisfies condition $(i)$ but not condition $(ii)$. It is also possible that two different UV Wilson loops, $W_{\vec{Q}_{1}}$ and $W_{\vec{Q}_{2}}$, map to the same simple object in the IR. In that case, we expect
$\begin{split}&\langle W^{+}_{\vec{Q}_{1}}W^{-}_{\vec{Q}_{2}}\rangle_{\rm sci}(q,\eta=1,\nu=-1)=\pm q^{\mathbb{Z}/2}\;.\end{split}$ (36)
In general, a UV line operator $L_{i}^{UV}$ that maps to an IR simple line may not correspond to any supersymmetric Wilson loop. Thus we expect that this procedure gives only a subset of the simple objects in the IR twisted theory.
@@ 3.4 Half-index and modular functions
Finally, we consider the half-index of the gauge theory decorated by the supersymmetric line operator $L_{i}^{UV}$, which is defined as Gadde:2013sca ; Gadde:2013wq ; Yoshida:2014ssa ; Dimofte:2017tpi
$I^{L_{i}}_{\rm half}(q,\nu,\eta):=\textrm{Tr}_{{\cal H}^{L_{i}}(D_{2};\mathbb{B})}(-1)^{R_{\nu}}q^{R_{\nu}/2+j_{3}}\eta^{A}.$ (37)
The trace is taken over the Hilbert space on the boundary torus with a specific boundary condition $\mathbb{B}$ and in the presence of the line operator $L_{i}^{UV}$.
In order to make contact with the Nahm sum formula, we impose Dirichlet boundary conditions ($\mathcal{D}$) for all ${\cal N}=2$ vector multiplets and deformed Dirichlet boundary conditions ($D_{c}$) for all ${\cal N}=2$ chiral multiplets. The half-index of ${\cal T}[K,\{{\cal O}_{I}\}]$ with the insertion of a supersymmetric Wilson loop $W_{\vec{Q}}$ then reads Dimofte:2017tpi
@@
with the Pochhammer symbol defined as $(q)_{n}:=\prod_{i=1}^{n}(1-q^{i})$. If this boundary condition is compatible with the topological A-twist in the infrared  [^10] This is in general a very non-trivial condition due to the existence of superpotentials, we can identify this specialization of the half-index with the character of the boundary VOA of the A-twisted theory:
$q^{\Delta}I_{A}[W_{\vec{Q}}]:=q^{\Delta}I^{W_{\vec{Q}}}_{\rm half}(q,\nu=-1,\eta=1)=\sum_{m\in\mathbb{N}^{r}}\frac{q^{\frac{1}{2}m^{t}Km+\Delta}(-q^{1/2})^{-(\mu_{0}-a)^{t}m}q^{-Q^{t}m}}{(q)_{m_{1}}\ldots(q)_{m_{r}}}\ .$ (39)
If ${\cal T}[K,\{{\cal O}_{I}\}]$ flows to a rank-zero SCFT and $W_{\vec{Q}}$ maps to a simple object in the IR $A$-twisted theory, it is reasonable to conjecture that (39) is a modular function with a proper choice of $\Delta\in\mathbb{Q}$. We can make a similar conjecture for the $B$-twist, and two sets $W^{\rm sim}_{A}$ and $W^{\rm sim}_{B}$ are related to each other by
$W^{\rm sim}_{B}=\{W_{(\vec{a}-\vec{Q})}\;:\;W_{\vec{Q}}\in W^{\rm sim}_{A}\}\;.$ (40)
Namely, the two sets of modular functions in A/B-twisting are identical.
For a half-index to have nice properties under modular transformations, it is natural to impose (NS, NS) boundary conditions on the boundary torus. The definition of the half-index (37) graded by $(-1)^{R}$ aligns with this choice. Further, the factor $(-1)^{(\mu_{0}-a)^{t}m}$ in the half-index, which does not appear in the original Nahm sum formula (3), also arises from this choice of spin structure. This implies we are working with a generalization of Nahm’s conjecture, as stated at the end of Section 2 and discussed in Appendix A.4. In Section 4, we will provide explicit examples of RCFT characters that can be expressed in this modified Nahm sum form.
@@ 4.1 A family of rank-zero theories
In this section, we summarize the classes of candidate rank-zero theories whose partition functions meet the criteria stated in Section 3. For each candidate, we also identify the set $W_{A}^{\rm sim}$ of UV Wilson loops which flow to simple objects in the IR A-twisted theory based on the criteria in Section 3.3. Through these examples, we confirm the conjecture in the previous section by expressing $I_{A}[W_{\vec{Q}}]$ in terms of the known characters of RCFTs. For the naming convention and explicit expressions of the characters, refer to Appendix B. The identification of the indices with known RCFT characters are conjectures based on the $q$-series expansion except for a few cases with known Nahm sum representations.
The search scanned all positive definite integer $K$-matrices for $r=1,2,3$ with entries between $-17$ and $17$. To avoid redundancy, we exclude candidates that are related to those with lower $r$ through basic mirror duality, or those which are constructed via direct products of other candidates. The basic mirror duality is expressed asAharony:1997bx ; Dimofte:2011ju
@@
Under this duality, a gauge-invariant monopole operator $V_{n\geq 0}$ is mapped to $\phi^{n}$. By deforming the two theories with a superpotential $V_{2}\simeq\phi^{2}$ (i.e. complex mass deformation) we find
$\left(\frac{{\cal T}_{\Delta}}{U(1)_{1}}\textrm{ with a superpotential }\mathcal{W}=V_{2}\right)\simeq(\textrm{an almost trivially gapped theory})\;.$ (42)
This implies that
@@
Here, $\tilde{{\cal O}_{I}}$ shares the same $(n_{i},m_{i})$ as ${\cal O}_{I}$ for $11\leq i\leq r-$ 1, but has $(n_{r},m_{r})=(0,0)$. At the level of the half-index, the duality can be understood from the following equality:
@@
Here $\chi_{F}$ is the character of the free Majorana fermion theory defined as
@@ 4.1.1 $r=1$
    $K=(2)=C(A_{1})$,
    $\begin{split}&\{{\cal O}_{I}\}=\emptyset,\quad\vec{\mu}_{0}=(-1),\quad\vec{a}=(1),\quad W_{\text{sim}}=\{1,W_{1}\},\\ &I_{\rm A}[1]=q^{-11/60}\chi^{M(2,5)}_{(1,1)},\quad I_{\rm A}[W_{1}]=q^{1/60}\chi^{M(2,5)}_{(1,2)},\\ &{\cal I}_{\rm sci}=1-q-\left(\eta+\frac{1}{\eta}\right)q^{3/2}-2q^{2}-\left(\eta+\frac{1}{\eta}\right)q^{5/2}-2q^{3}-\left(\eta+\frac{1}{\eta}\right)q^{7/2}+\cdots\,.\\ \end{split}$
@@ 4.1.2 $r=2$
In addition to the examples which are direct products of two $r=1$ theories, we have three candidate rank-zero theories.
@@
    This example is studied in Baek:2024tuo . In particular, it was argued that the supersymmetry at the fixed point enhances to ${\cal N}=5$. Based on various partition function calculations, it is conjectured that this is dual to ${\cal N}=3$ $SU(2)_{k=2}$ Chern-Simons theory coupled to a half hypermultiplet and a half twisted hypermultiplet both in the fundamental representation.
@@
    $\begin{split}&\{{\cal O}_{I}\}=\{V_{(-1,2)}\},\quad\vec{\mu}_{0}=(-2,-1),\quad\vec{a}=(2,1),\quad W^{\rm sim}_{A}=\{1,W_{(1,1)},W_{(2,1)}\},\\ &I_{\rm A}[1]=q^{-17/42}\chi^{M(2,7)}_{(1,1)},\quad I_{\rm A}[W_{(1,1)}]=q^{-5/42}\chi^{M(2,7)}_{(1,2)},\quad I_{\rm A}[W_{(2,1)}]=q^{1/42}\chi^{M(2,7)}_{(1,3)},\\ &{\cal I}_{\rm sci}=1-q-\left(\eta+\frac{1}{\eta}\right)q^{3/2}-2q^{2}-\eta q^{5/2}+\left(\frac{1}{\eta^{2}}-1\right)q^{3}+\left(\frac{1}{\eta}-\eta\right)q^{7/2}+\frac{1}{\eta^{2}}q^{4}+\cdots.\\ \end{split}$
    This example is discussed in Gang:2023rei ; Ferrari:2023fez ; Creutzig:2024ljv in detail, where it is argued that the boundary condition indeed supports the $M(2,7)$ Virasoro minimal model.
@@ 4.1.3 $r=3$
In addition to the examples which are direct products of $r=1$ and $r=2$ theories, we find 23 candidate rank-zero theories. Many of these theories have identical superconformal indices and partition functions, which is a strong signal that they flow to the same fixed point. Below we organize the theories by duality class.
@@
The following list of theories have the superconformal index
@@
which is the same as that of ${\cal T}_{\text{min}}$. There exists a gauge theory description whose (deformed) Dirichlet half-index reproduces the characters of $M(2,5)$ or the simple affine VOA $L_{1}(\mathfrak{osp}(1|2))$ (with a specialization of the Jacobi variable) up to a free fermion factor $\chi_{F}$.
@@
    $\begin{split}&\{{\cal O}_{I}\}=\{(\phi_{1})^{2}V_{(0,0,2)},(\phi_{3})V_{(1,1,0)}\},\quad\vec{\mu}_{0}=(-1,1,0),\quad\vec{a}=(1,-1,0),\quad W^{\rm sim}_{A}=\{1,W_{(1,-1,0)}\},\\ &I_{\rm A}[1]=q^{3/80}\chi_{F}(q)\chi^{osp(1|2)_{1}}\left[{\bf 1}\right](q,x)|_{x=-q^{1/2}},\\ &I_{\rm A}[W_{(1,-1,0)}]=q^{27/80}\chi_{F}(q)\chi^{osp(1|2)_{1}}\left[{\bf M}\right](q,x)|_{x=-q^{1/2}}.\end{split}$
    Here $\chi^{osp(1|2)_{1}}[{\bf 1}]$ and $\chi^{osp(1|2)_{1}}[{\bf M}]$ are the supercharacters of the vacuum module 1 and a non-vacuum module ${\bf M}$ of $L_{1}(\mathfrak{osp}(1|2))$. (See appendix B for the notation.)
@@
    $K=\begin{pmatrix}2&-1&-1\\ -1&2&1\\ -1&1&1\\ \end{pmatrix}\ ,$
@@
    $K=\begin{pmatrix}2&1&-1\\ 1&2&0\\ -1&0&1\\ \end{pmatrix}\ ,$
@@
    $K=\begin{pmatrix}2&1&1\\ 1&2&1\\ 1&1&1\\ \end{pmatrix},\$
@@
This theory is expected to flow to two copies of ${\cal T}_{\text{min}}$. As evidence, we can check that the superconformal index is a square of (46):
@@
The following class of theories has the superconformal index
which is the index of ${\cal T}_{\text{min}}$ multiplied by that of $U(1)_{2}$ theory.
@@
    $K=\begin{pmatrix}4&-2&2\\ -2&2&-1\\ 2&-1&2\end{pmatrix},$  [^12] For the second example, we find ${\cal Z}^{\text{top}}_{S^{2}\times S^{1}}=0$, although the superconformal index is equal to 1 at the A/B-twisted points. See section 5 for more discussion of this example.
@@
The following class of theories has the superconformal index
@@
which is the same as the rank-zero theory ${\cal T}_{2}$ introduced in Gang:2023rei ; Ferrari:2023fez . This is also the superconformal index of the theories ${\cal T}_{1,2}(={\cal T}_{2})$ and $\overline{{\cal T}}_{2,1}$ in Creutzig:2024ljv . The latter two theories are related by mirror symmetry
@@
where ${\cal T}^{\vee}$ and $\overline{{\cal T}}$ denote the mirror dual and the orientation reversal of ${\cal T}$ respectively. It is argued in loc. cit. that the Dirichlet half-index of the A- and B-twist of ${\cal T}_{1,2}$ (with a suitable specialization) gives the characters of $L_{2}(\mathfrak{osp}(1|2))$ and $M(2,7)$ respectively. The Dirichlet half-index of the A- and B-twist of ${\cal T}_{2,1}$ reproduces the characters of $L_{1}(\mathfrak{osp}(1|4))$ and the minimal W-algebra $W^{\text{min}}_{1/2}(\mathfrak{sp}(4))$, which are also realized at the left boundary of the mirror dual theory ${\cal T}_{1,2}$.
@@
The following theory has the superconformal index
@@
The following theory has the superconformal index
@@
which agrees with the superconformal index of the theory ${\cal T}_{(2,8)}$ in Baek:2024tuo . The Dirichlet half-index of this theory reproduces the vacuum character of the ${\cal N}=1$ super Virasoro minimal model $SM(2,8)$.
@@
The following class of theories has the superconformal index
@@
which agrees with the superconformal index of the theory ${\cal T}_{(2,12)}$ defined in Baek:2024tuo . The Dirichlet half-index of the theory reproduces the vacuum character of the ${\cal N}=1$ super Virasoro minimal model $SM(2,12)$.
@@
The following theory has the superconformal index
@@
$-(\eta+1/\eta)q^{3/2}$ (56)
in the $q$-expansion. The existence of this term is a strong signal that the supersymmetry is enhanced to ${\cal N}=4$, since it coincides with the contribution from the extra supercurrent multiplet. However, it is neither a necessary nor sufficient condition, as the spectrum may contain other multiplets whose contribution to the index is the same, possibly with the opposite sign.

### v2

@@ 3.3 Simple lines
Suppose that an ${\cal N}=2$ gauge theory $T[K,\{{\cal O}_{I}\}]$ flows to an ${\cal N}=4$ rank-zero SCFT. Then there exists a distinguished set of line operators $\{L_{1}^{\text{UV}},\cdots L_{d}^{\text{UV}}\}$ in the UV gauge theory which flow to simple lines (i.e. anyons) in the IR topologically twisted field theory. In this section, we focus on the case where UV line operators take the form of supersymmetric Wilson loops $W_{\vec{Q}}$ with charge $\vec{Q}=(Q_{1},\ldots,Q_{r})$, and denote $W^{\rm sim}_{A/B}$ by the collection of loops which map to simple lines. The two sets $W^{\rm sim}_{A}$ and $W^{\rm sim}_{B}$ always contain the trivial Wilson line, $1=W_{\vec{Q}=\vec{0}}$, which corresponds to the trivial simple line in the twisted TFT. Below we list several necessary conditions for a UV operator $W_{\vec{Q}}$ to become a simple object in the IR.
@@
Let us consider the $S^{2}\times S^{1}$ partition function of the $A$-twisted theory, which can be computed by the superconformal index with specialization $\nu=-1$, as discussed in section 3.2. In order to test whether a UV line operator $W_{\vec{Q}}$ flows to a simple line in the IR A-twisted TFT, we scan over integer charge vectors $\vec{Q}$ and search for $W_{\vec{Q}}$ which satisfy the following two necessary conditions  [^9] More precisely, the conditions for $W_{\vec{Q}}$ to become a simple object, up to a factor of $U(1)_{R}$ flavor Wilson loop which contributes $(-q^{\pm 1/2})^{Q_{R}\in\mathbb{Z}}$ to $\langle W_{\vec{Q}}^{\pm}\rangle_{\rm sci}$. These contributions cancel in $\langle W^{+}_{\vec{Q}}W^{-}_{\vec{Q}}\rangle_{\rm sci}$.
@@
The first condition follows from the fact that the $S^{2}\times S^{1}$ partition function for a semi-simple TFT with a non-trivial simple line insertion either vanishes or is equal to 1 if we insert the identity line. The second condition follows from the fact that the fusion between a simple object $\mathbf{L}$ and its orientation reversal $\overline{\mathbf{L}}$ always contains a single $1$ (trivial object), i.e.,
$\mathbf{L}\times\overline{\mathbf{L}}\sim 1+(\textrm{non-trivial simple objects})\;.$ (39)
More concretely, we have
@@
Here $S_{\alpha\beta}$ is the modular S-matrix and $(S^{2})_{\alpha\beta}=C_{\alpha\beta}$ is the charge-conjugation matrix. If a UV Wilson loop maps to a sum of multiple simple objects in the IR, it satisfies condition $(i)$ but not condition $(ii)$. It is also possible that two different UV Wilson loops, $W_{\vec{Q}_{1}}$ and $W_{\vec{Q}_{2}}$, map to the same simple object in the IR. In that case, we expect
[…]
In general, a UV line operator $L_{i}^{UV}$ that maps to a IR simple line may not correspond to any supersymmetric Wilson loop. Thus we expect that this procedure gives only a subset of the simple objects in the IR twisted theory.
@@ 3.4 Half-index and modular functions
Finally, we consider an ${\cal N}=2$ gauge theory $T[K,\{{\cal O}_{I}\}]$ defined on the Euclidean half-space $\mathbb{C}\times\mathbb{R}_{\leq 0}$. The local operators on the boundary are counted by the half-index, which can be understood as the partition function on the cigar geometry $D^{2}\times S^{1}$, with a choice of boundary condition $\mathbb{B}$ specified on the boundary torus. Gadde:2013sca ; Gadde:2013wq ; Yoshida:2014ssa ; Dimofte:2017tpi Let $(z,\bar{z},t)$ be the local coordinate around a point on the boundary placed at $t=0$. The half index is defined by
$I_{\rm half}(q,\nu,\eta):=\textrm{Tr}_{T^{2}}(-1)^{R_{\nu}}q^{R_{\nu}/2+j_{3}}\eta^{A}\ .$ (42)
It can be further decorated by a loop operator $L$ placed at the tip of the cigar, wrapping $S^{1}$. We denote this by $I^{L}_{\rm half}(q,\nu,\eta)$.
In order to make contact with the Nahm sum formula discussed in Section 2, we impose the supersymmetric Dirichlet boundary conditions that preserve a 2d ${\cal N}=(0,2)$ subalgebra, as discussed in Dimofte:2017tpi . Concretely, for an ${\cal N}=2$ vector multiplet, we take
$A_{z}|_{t=0}=A_{\bar{z}}|_{t=0}=0\ ,\quad D|_{t=0}=0\ ,\quad\lambda_{-}|_{t=0}=0\ ,$ (43)
which we denote by the boundary condition $\cal D$}. Crucially, this boundary condition admits the boundary monopole operators, which are solutions to the BPS equations
$F=*D\sigma\ ,\quad\sigma=\frac{1}{(|z|^{2}+t^{2})^{1/2}}\ ,$ (44)
labeled by the monopole charge
$\int_{D^{2}}\frac{F}{2\pi}=m\in\mathbb{Z}^{r}\ .$ (45)
For an ${\cal N}=2$ chiral multiplet, we impose
$\phi|_{t=0}=c\ ,\quad\psi_{+}|_{t=0}=0\ ,$ (46)
where $c\neq 0$ is a complex number. We denote this boundary condition by $D_{c}$, a deformed Dirichlet boundary condition. The half-index of ${\cal T}[K,\{{\cal O}_{I}\}]$ with the insertion of a supersymmetric Wilson loop $W_{\vec{Q}}$ then reads
@@
with the Pochhammer symbol $(q)_{n}:=\prod_{i=1}^{n}(1-q^{i})$ defined in the usual way. The summation over $m$ labels the boundary monopole operators discussed above.
Similarly to the discussion for the superconformal index, the half index specialized at $\nu=0$ calculates the character of the vector space of boundary local operators at the superconformal fixed point, with the R-charge determined by F-maximization. Now, let us assume that the supersymmetry at the fixed point enhances to ${\cal N}=4$, and moreover that the boundary condition $\mathbb{B}$ in the infrared is compatible with the supercharges of the twisted theory $Q_{A}$ or $Q_{B}$. Then one can define the half-index of the topologically twisted theories, which coincides with (42) specialized at $\nu=\pm 1$ and $\eta=0$,
$I_{\text{half}}(q,1,-1)=~\textrm{Tr}_{T^{2}}(-1)^{R_{\nu}}q^{J_{3}^{H}+j_{3}}\ ,\quad I_{\text{half}}(q,1,1)=~\textrm{Tr}_{T^{2}}(-1)^{R_{\nu}}q^{J_{3}^{C}+j_{3}}\ .$ (48)
Therefore, at $\nu=-1$, we expect that the half index calculates the vacuum character of the corresponding boundary vertex operator algebra of the A-twisted theory,
$\chi_{A}(q)=q^{\Delta}I_{\rm half}(q,\nu=-1,\eta=1)=\sum_{m\in\mathbb{N}^{r}}\frac{q^{\frac{1}{2}m^{t}Km+\Delta}(-q^{1/2})^{-(\mu_{0}-a)^{t}m}}{(q)_{m_{1}}\ldots(q)_{m_{r}}}\ .$ (49)
The non-vacuum characters for the simple modules can be computed by inserting the line operator that flows to a simple line,
$q^{\Delta}I^{W_{\vec{Q}}}_{\rm half}(q,\nu=-1,\eta=1)=\sum_{m\in\mathbb{N}^{r}}\frac{q^{\frac{1}{2}m^{t}Km+\Delta}(-q^{1/2})^{-(\mu_{0}-a)^{t}m}q^{-Q^{t}m}}{(q)_{m_{1}}\ldots(q)_{m_{r}}}\ ,$ (50)
which we denote by $q^{\Delta}I_{A}[W_{\vec{Q}}]$.
Notice that the assumption that the boundary condition $\mathbb{B}$ is compatible with the supercharges $Q_{A}$ or $Q_{B}$ is highly non-trivial, due to the presence of superpotentials involving monopole operators.  [^10] See section 5 of Costello:2020ndc for discussion on the classification of boundary conditions in the presence of superpotential deformation. For a special class of abelian Chern-Simons matter theories studied in the work of the second author Ferrari:2023fez ; Creutzig:2024ljv , there is strong evidence that this condition can indeed be satisfied. In this paper, we do not attempt a detailed analysis of the $Q_{A/B}$-invariance of the boundary condition $\mathbb{B}$ in the IR topological theory. Instead, we focus on analzing the expressions (50) for the theories ${\cal T}[K,\{{\cal O}_{I}\}]$ that are expected to flow to rank-zero SCFTs, and propose candidate boundary VOAs supported by the corresponding TFTs when the conditions are satisfied. We indeed find in section 4 that there are examples (see e.g., 3-6, 3-7 and 3-8) where the index identically vanishes, suggesting that the the boundary condition $\mathbb{B}$ may not preserve the twisted supercharges.
If ${\cal T}[K,\{{\cal O}_{I}\}]$ flows to a rank-zero SCFT and $W_{\vec{Q}}$ maps to a simple object in the IR $A$-twisted theory, we expect that (50) is a modular function under a congruence subgroup of $SL(2,\mathbb{Z})$, with a proper choice of $\Delta\in\mathbb{Q}$. We can make a similar conjecture for the $B$-twist.
For a half-index to have nice properties under modular transformations, it is natural to impose the (NS, NS) boundary condition on the boundary torus. The definition of the half-index (42) graded by $(-1)^{R}$ aligns with this choice. Further, the factor $(-1)^{(\mu_{0}-a)^{t}m}$ in the half-index, which does not appear in the original Nahm sum formula (3), arises from this choice of spin structure. This implies we are working with a generalization of Nahm’s conjecture, as stated at the end of Section 2 and discussed in Appendix A.4. In Section 4, we will provide explicit examples of RCFT characters that can be expressed in this modified Nahm sum formula.
@@ 4.1 A family of rank-zero theories
In this section, we summarize the class of the candidate rank-zero theories whose partition functions meet the conditions stated in section 3. For each candidate, we also identify the set $W_{A}^{\rm sim}$ of UV Wilson loops that flow to simple objects in the IR A-twisted theory, based on the criteria in section 3.3. Through these examples, we confirm the conjecture in the previous section by expressing the $I_{A}[W_{\vec{Q}}]$ in terms of the known characters of RCFTs. For the naming convention and explicit expressions of the characters, refer to Appendix B. The identification of the indices with known RCFT characters are conjectures based on the $q$-series expansion, except for a few cases with known Nahm sum representations.
The search scanned all positive definite integer $K$-matrices for $r=1,2,3$ with entries between 17-17 and $17$. To avoid redundancy, we exclude candidates that are related to those with lower $r$ through basic mirror duality or direct products of lower $r$ candidates. The basic mirror duality is expressed as Aharony:1997bx ; Dimofte:2011ju
@@
Under this duality, a gauge-invariant monopole operator $V_{n\geq 0}$ is mapped to $\phi^{n}$. By deforming the two theories with a superpotential $V_{2}\simeq\phi^{2}$, i.e. complex mass deformation, we find
[…]
It implies that
@@
Here, $\tilde{{\cal O}_{I}}$ shares the same $(n_{i},m_{i})$ as ${\cal O}_{I}$ for $1\leq i\leq r-1$, but has $(n_{r},m_{r})=(0,0)$. At the level of the half-index, the duality can be understood from the following equality
@@
Here $\chi_{F}$ is the character of the free Majorana fermion theory,
@@ 4.1.1 $r=1$
    $K=(2)=C(A_{1})$, 121212 Taking into account the background CS level 1/2-1/2 in the ${\cal T}_{\Delta}$ theory (2), this theory corresponds to the $U(1)_{3/2}$ theory coupled to a single chiral multiplet of charge +1+1, which we denote simply as $U(1){3/2}+\Phi$. This theory is IR dual to $U(1)_{-3/2}+\Phi$ Gang:2018huc .
[…]
@@ 4.1.2 $r=2$
Besides the examples which are direct products of two $r=1$ theories,
@@
    This example is studied in Baek:2024tuo . In particular, it was argued that the supersymmetry at the fixed point enhances to ${\cal N}=5$. Based on various partition function calculations, it is conjectured that this is dual to the ${\cal N}=3$ $SU(2)_{k=2}$ Chern-Simons theory coupled to a half hypermultiplet and a half twisted-hypermultiplet both in fundamental representation.
@@
[…]
    This example is discussed in Gang:2023rei ; Ferrari:2023fez ; Creutzig:2024ljv in detail, where it is argued that the boundary condition indeed supports the $M(2,7)$.
@@ 4.1.3 $r=3$
Besides the examples which are direct products of $r=1$ and $r=2$ theories, we find 23 candidate rank-zero theories. Many of these theories have identical superconformal indices and partition functions, which is a strong signal that they flow to the same fixed point. Below we organize the theories by the duality class.
@@
The following list of the theories have the superconformal index
@@
which is the same as that of ${\cal T}_{\text{min}}$. There exists a gauge theory description whose (deformed) Dirichlet half-index reproduces the characters of $M(2,5)$ or the simple affine VOA $L_{1}(\mathfrak{osp}(1|2))$ (with a specialization of the Jacobi variable), up to a free fermion factor $\chi_{F}$.
@@
    $\begin{split}&\{{\cal O}_{I}\}=\{(\phi_{1})^{2}V_{(0,0,2)},(\phi_{3})V_{(1,1,0)}\},\quad\vec{\mu}_{0}=(-1,1,0),\quad\vec{a}=(1,-1,0),\quad W^{\rm sim}_{A}=\{1,W_{(1,-1,0)}\},\\ &I_{\rm A}[1]=q^{3/80}\chi_{F}(q)\chi^{osp(1|2)_{1}}\left[{\bf 1}\right](q,x)|_{x=-q^{1/2}},\\ &I_{\rm A}[W_{(1,-1,0)}]=q^{27/80}\chi_{F}(q)\chi^{osp(1|2)_{1}}\left[{\bf M}\right](q,x)|_{x=-q^{1/2}},\end{split}$
    where $\chi^{osp(1|2)_{1}}[{\bf 1}]$ and $\chi^{osp(1|2)_{1}}[{\bf M}]$ are the supercharacters of the vacuum module 1 and a non-vacuum module $\bf M$} of $L_{1}(\mathfrak{osp}(1|2))$. (See appendix B for the notation.)
@@
    $K=\begin{pmatrix}2&-1&-1\\ -1&2&1\\ -1&1&1\\ \end{pmatrix}\$,
@@
    $K=\begin{pmatrix}2&1&-1\\ 1&2&0\\ -1&0&1\\ \end{pmatrix}\$,
@@
    $K=\begin{pmatrix}2&1&1\\ 1&2&1\\ 1&1&1\\ \end{pmatrix}$,\\
@@
This theory is expected to flow to the two copies of ${\cal T}_{\text{min}}$’s. As the first evidence, we can check that the superconformal index is a square of (56):
@@
${\cal I}_{\rm sci}=1-q-\left(\eta+\frac{1}{\eta}\right)q^{3/2}-2q^{2}-\left(\eta+\frac{1}{\eta}\right)q^{5/2}-2q^{3}-\left(\eta+\frac{1}{\eta}\right)q^{7/2}+\cdots\ .$ (59)
@@
    $K=\begin{pmatrix}4&-2&2\\ -2&2&-1\\ 2&-1&2\end{pmatrix}$,  [^13] For the second example, we find $Z^{\text{top}}_{S^{2}\times S^{1}}=0$, although the superconformal index is equal to 1 at the A/B-twisted points. See section 5 for more discussion on this example.
@@
The following class of the theories have the superconformal index expansion
@@
which is the same as the rank-zero theory ${\cal T}_{2}$ introduced in Gang:2023rei ; Ferrari:2023fez . This is also the superconformal index of the theories ${\cal T}_{1,2}(={\cal T}_{2})$ and $\overline{{\cal T}}_{2,1}$ in Creutzig:2024ljv . The latter two theories are related by the mirror symmetry:
@@
where $\cal T$$^{\vee}$ and $\overline{{\cal T}}$ denotes the mirror dual and the orientation reversal of $\cal T$} respectively. It is argued in loc. cit. that the Dirichlet half-index of the A- and B-twist of ${\cal T}_{1,2}$ (with a suitable specialization) gives the characters of $L_{2}(\mathfrak{osp}(1|2))$ and $M(2,7)$ respectively. The Dirichlet half-index of the A- and B-twist of ${\cal T}_{2,1}$ reproduces the characters of $L_{1}(\mathfrak{osp}(1|4))$ and the minimal W-algebra $W^{\text{min}}_{1/2}(\mathfrak{sp}(4))$, which are also realized at the left boundary of the mirror dual theory ${\cal T}_{1,2}$.
@@
The following theory has the superconformal index expansion
@@
The following class of theories have the superconformal index expansion:
@@
which agrees with the superconformal index of the theory ${\cal T}_{(2,8)}$ in Baek:2024tuo . The Dirichlet half-index of the theory reproduces the vacuum character of $N=1$ super Virasoro minimal model $SM(2,8)$.
@@
The following class of theories have the superconformal index expansion:
@@
which agrees with the superconformal index of the theory ${\cal T}_{(2,12)}$ defined in Baek:2024tuo . The Dirichlet half-index of the theory reproduces the vacuum character of $N=1$ super Virasoro minimal model $SM(2,12)$.
@@
The following theory has the superconformal index expansion
@@
$(\eta+1/\eta)q^{3/2}$ (66)
in the $q$-expansion. The existence of this term is a strong signal that the supersymmetry is enhanced to ${\cal N}=4$, since it coincides with the contribution from the extra supercurrent multiplet. However, it is neither a necessary nor a sufficient condition, as the spectrum may contain other multiplets whose contribution to the index is the same, possibly with the opposite sign.